Financial information for Powell Panther Corporation is shown below: Powell Panther Corporation: Income Statements for Year Ending December 31 (Millions of Dollars) 2016 2015 $3,375.0 $2,700.0 2.616.0 2.295.0 $759.0 $405.0 Sales EBITDA $688.0 $346.0 24.0 59.0 $614.0 $287.0 245.6 114.8 $368.4 $172.2 Earnings before interest and taxes (EBIT) Interest Earnings before taxes (EBT) Taxes (40%) Net income Common dividends $332.0 $138.0 Powell Panther Corporation: Balance Sheets as of December 31 (Millions of Dollars) 2016 2015 Assets Cash and equivalents Accounts receivable $37.0 $32.0 456.0351.0 944.0 575.0 $1.337.0 $1,058.0 213.0 94.0 $2,050.0 $1,652.0 Total current assets Net plant and equipment Total assets Liabilities and Equity Accounts payable $351.0 $270.0 119.0 108.0 68.0 54.0 $538.0 $432.0 675.0 540.0 $1.213.0$972.0 749.6 629.0 Notes payable Total current liabilities Long-term bonds Total liabilities Common stock Retained earnings $830 $6800 $2.050.0 $1,652.0 Common equity Total liabilities and equity Write out your ansvers completely, For example, 25 million should be entered as 25.000.000. a. What was net operating working capital for 2015 and 2016?
Write out your answers completely. For example, 25 million should be entered as 25,000,000. a. What was net operating working capital for 2015 and 2016? 2015: $ 2016: $ b. What was the 2016 free cash flow? c. How would you explain the large increase in 2016 dividends? I. The large increase in net income from 2015 to 2016 explains the large increase in 2016 dividends. II. The large increase in EBIT from 2015 to 2016 explains the large increase in 2016 dividends. II. The large increase in free cash flow from 2015 to 2016 explains the large increase in 2016 dividends. The large increase in sales from 2015 to 2016 explains the large increase in 2016 dividends. V. The large increase in retained earnings from 2015 to 2016 explains the large increase in 2016 dividends. -Select-
Answer #1

1. Net Operating Working Capital (NOWC) = Net Operating Assets-Net Operating Liabilities

NOWC = Cash + Accounts Receivable + Inventories - Accounts Payable - Accrued Expenses

NOWC for 2015 = (32 + 351 + 675 - 270 - 108) in million dollars

NOWC for 2015 = $680,000,000

Applying same formula for NOWC 2016, it gives -

NOWC for 2016 = (37 + 456 + 844 - 351 - 119) in million dollars

NOWC for 2016 = $867,000,000

2. Free Cash Flow (FCF) = EBIT * (1-t) +/- changes in Working Capital - Capital Expenditure

FCF for 2016 = 688 (1-0.4) + 71 + 799 - 119

FCF for 2016 = $1,083,800,000

3. Option 1 is correct because increase in dividends in majorly caused primarily by 2 reasons:

a. Increase in net income of the company,

b. A shift in the company’s growth strategy that leads the company to decide to expend less of its earnings in seeking growth and expansion, thus leaving a larger share of profits available to be returned to equity investors in the form of dividends.

Since, we do not have any information regarding the second reason, we can consider that dividend increase has been caused by increase in net income of the company.